OpenStudy (anonymous):

log_5(x^2)-log_5(x+5)=6 x=?

OpenStudy (watchmath):

\(\log_5(\frac{x^2}{x+5})=6\) \(5^6=\frac{x^2}{x+5}\) \(3125(x+5)=x^2\) \(x^2-3125x-5\cdot 3125=0\) Then solve by quadratic formula :)

OpenStudy (anonymous):

got it thank you!!!